# Deriving Representative Structure from Music Corpora

### v2

@@ Abstract
Western music is an innately hierarchical system of interacting levels of structure, from fine-grained melody to high-level form. In order to analyze music compositions holistically and at multiple granularities, we propose a unified, hierarchical meta-representation of musical structure called the structural temporal graph (STG). For a single piece, the STG is a data structure that defines a hierarchy of progressively finer structural musical features and the temporal relationships between them. We use the STG to enable a novel approach for deriving a representative structural summary of a music corpus, which we formalize as a dually NP-hard combinatorial optimization problem extending the Generalized Median Graph problem. Our approach first applies simulated annealing to develop a measure of structural distance between two music pieces rooted in graph isomorphism. Our approach then combines the formal guarantees of SMT solvers with nested simulated annealing over structural distances to produce a structurally sound, representative centroid STG for an entire corpus of STGs from individual pieces. To evaluate our approach, we conduct experiments verifying that structural distance accurately differentiates between music pieces, and that derived centroids accurately structurally characterize their corpora.
@@ 1 Introduction
A prevailing theory among Western music theorists and musicologists states that Western classical music exhibits an implicitly hierarchical structure [?]. While several different theoretical systems have been proposed to formalize this structural hierarchy [?], a widely accepted modern interpretation of the hierarchy states that melodies form the bottom, followed by harmonic contour, rhythmic patterns, disjoint and possibly overlapping motifs, and finally large-scale sections [???]. Together, this composite hierarchy encapsulates the overall structure of a piece.
To analyze musical structure computationally, many automatic approaches have been developed for extracting structure at single levels of the structural hierarchy [?????]. However, music perception researchers have shown that the levels are not perceptually independent: they relate to one another in both “vertical” (structural) and “horizontal” (temporal) directions [?], interactions that have more recently been proven computationally  [?]. A comprehensive analysis of a piece thus must integrate these inter- and intra-level interactions into a unified model. Furthermore, while attempts have been made to generate structured music [???], to our knowledge, no existing research has examined generating music adhering to a complete, musically representative structural hierarchy, since no mechanism to computationally derive a complete structural specification from a desired music corpus currently exists. Such an encapsulation of music structure could also play a critical role in the generation of well-formed music by serving as a system of constraints on generative models.
@@
To address these challenges, we introduce the structural temporal graph (STG) as a unified model of complete musical structure. The STG is a $k$-partite directed acyclic graph whose levels form the structural hierarchy, and edges encode temporal relationships between adjacent levels. We use simulated annealing to develop a measure of structural distance between two STGs based on graph isomorphism, and to obtain the overarching structure of a corpus of pieces, we develop an approach to derive a representative centroid graph from a corpus of STGs. We formalize centroid derivation as a dually NP-hard combinatorial optimization problem extending the Generalized Median Graph problem [?], and propose a solution combining nested simulated annealing with the formal guarantees of SMT solvers to produce a structurally sound result. Our experiments show that structural distance accurately differentiates pieces, with its performance reliant on the complete hierarchy, and that derived centroids accurately structurally characterize their corpora. [^1] All code can be accessed in our GitHub repository: github.com/ilanashapiro/constraints_project
@@
    We formalize the music summarization problem as a dually NP-hard combinatorial optimization problem, and contribute a novel solution using both stochastic and SMT-based techniques.
@@ 3 Structural Temporal Graph
To address the lack of a fully automatic complete encapsulation of polyphonic musical structure, we introduce the structural temporal graph (STG), a unified meta-representation of musical structure that captures the levels of the music structural hierarchy and the temporal relationships between them. The STG is a $k$-partite directed acyclic graph (DAG), where each of the $k$ layers encodes a level in the music structural hierarchy. [^2] Individual levels themselves can form sub-hierarchies of increasing granularity, which the STG supports. Following the modern music theoretic interpretation of the hierarchy [???], from top to bottom we denote the levels to be contiguous segmentation, motifs (both disjoint and overlapping), rhythmic contour, relative keys, functional harmonic chords, and melodic contour. The STGs we build include every level in the hierarchy except rhythmic contour, for which we were unable to access an analysis algorithm. We run individual analysis algorithms to generate each level of the hierarchy, which is elucidated in Section 6. Before formally defining the STG, we build intuition by walking through the derivation of an STG from an annotated piece.
@@ 4.1 Graph Alignment Annealing
GED is a NP-hard combinatorial optimization problem [?], intractable for most STGs. Most GED approximation algorithms are slow and of dubious accuracy, and more generalized than we require. We thus introduce new measure of structural distance computed with simulated annealing (SA). SA is a stochastic optimization technique used to estimate the global optimum of a discrete cost function. It comprises an objective “energy” function to minimize and a “move” function for generating a new solution from the current state. An annealer begins at a high temperature indicating the likelihood of accepting worse solutions to explore the solution space, and ends at a low temperature to refine near-optimal solutions  [?].
@@ 5.1 Bi-Level Simulated Annealing
The centroid annealer’s energy function is given in Equation 5, where $C_{\text{aligned}}$ is the corpus aligned to the current centroid $A_{g}$, a process itself requiring SA as in Section 4.1 to obtain the optimal alignments. Finding the centroid $A_{g}$ minimizing Equation 4 is thus a dually NP-hard problem (GED and minimizing over these distances) requiring nested SA.
@@
As the centroid annealer’s temperature cools, the loss converges as the centroid is increasingly closely aligned to its corpus. Thus, as the centroid annealer’s temperature cools, we scale down the number of steps and max temperature of the nested alignment annealer (precise cooling schedule is in Appendix Section A).
@@ 1. Every instance node must have exactly one prototype parent per feature
We first partition $A_{g}$ into pairs of consecutive instance levels without their prototypes (e.g. a segmentation/motif pair of instance levels), and optimize the instance constraints in Table 2 and relevant global constraints in Table 1 over each partition incrementally. We combine the results of each partition until we build a valid centroid subgraph of instance nodes. Then, we partition $A_{g}$ into single levels, each containing the instance nodes of that level and all prototype nodes for each instance feature at that level (e.g. segmentation instance nodes + section number prototype nodes). The instance constraints are already optimized; we need only optimize the prototype constraints in Table 2 and relevant global constraints in Table 1 over the possible prototypes. This gives us the complete, structurally sound centroid $A^{\prime}_{g}$ we seek. A detailed visualization of the complete centroid derivation process is in Appendix Section B.
@@ 6 Experiments
To evaluate our approach, we create a dataset of polyphonic, symbolic MIDI piano music from the Kunstderfuge and Classical Piano MIDI datasets 555 https://www.kunstderfuge.com/ and http://www.piano-midi.de/. Since some single-level analyses we use to generate STGs require the data to be in audio and CSV format, we convert MIDI to CSV (with a manual script) and to MP3 (with Fluidsynth [?]).
@@ 6.1 Structural Distance Evaluation
Musical Evaluation. To verify structural distance accurately differentiates pieces, we construct 210 STG sets from 32 pieces by composers J.S. Bach (21), Mozart (2), Beethoven (3), Schubert (2), and Chopin (4). Each set is a unique combination of 5 pieces, one from each composer, such that the duration of any piece is within 7 seconds of any other piece in the same set, since structural distance between disparate length pieces could be due to STG size differences rather than the local structural variations we aim to distinguish. Input pieces/set combination details are in Appendix Section C. We compute pairwise structural distances between the STGs in each set with our Section 4 graph alignment annealer (2000 iterations, max and min temperature of 2 and 0.01) on 8 Nvidia RTX 2080 GPUs with 11GB RAM. This results in 210 structural distance matrices, which we average to a single mean distance matrix.
We evaluate structural distance against three baselines over the same 210 piece combinations. Baseline 1 is the mean distance matrix obtained by taking the cosine similarity between feature vectors extracted from each MIDI file using Music21 [?]. Baseline 2 is the mean distance matrix over pairwise Stent weighted audio similarities (SWAS) for each piece’s paired MP3 file. SWAS is a composite audio similarity metric comprising zero-crossing rate, rhythm, chroma, spectral contrast, and perceptual similarity metrics, which we weigh equally. To demonstrate that existing graph comparison metrics are insufficient, Baseline 3 is the Weisfeiler-Lehman (WL) Kernel applied pairwise to the STGs in each set, with five iterations and Neighborhood Hash (NH) as the base kernel. The WL Kernel iteratively refines node labels based on their neighbors, and uses a base kernel (in this case, NH, which efficiently captures local graph structure by distinguishing neighborhood configurations) to compare them [?]. Our ground-truth is the stylistic similarity indices between composers using human annotations and metadata from “The Classical Music Navigator” [?].
@@ 6.2 Centroid Evaluation
Mathematical Verification. As a correctness check, we evaluate derived centroids against constructed ground truth centroids. Given base STG $g$ with $|E|$ edges, we create a synthetic corpus $C_{k}$ of $k$ STGs by randomly adding $\lceil\frac{|E|}{2}\rceil$ valid edits to $g$, $k$ times, as in Section 6.1. By construction, $g$ must be a true centroid with min possible loss over $C_{k}$ (Equation 4; proof in Appendix Section D). We derive candidate centroid $g_{d}$ from $C_{k}$, and compare to $g$. We evaluate $g_{d}$ against the naive centroid $g_{n}$, i.e., the STG already in $C_{k}$ with min loss over the rest of $C_{k}$.
@@
Musical Evaluation. This experiment requires pieces with more length (and STG size) variation than Section 6.1, as centroids derived from very similar input graphs may be trivial by construction and fail to generalize to more diverse corpora. We thus relax the Section 6.1 relative duration restriction to 80 seconds, and generate centroid STGs for four corpora with pieces by Alkan (11), Chopin (8), Haydn (12), and Mozart (14) with the same architecture as before. Input piece details are in Appendix Section C.
To verify each derived centroid $g$ musically characterizes its corpus $C$, we use rustworkx [?] to enumerate the set of all 5-node subgraphs ($\{S_{5}\}$) common to every STG in $C$, thus extracting the most structurally salient musical relationships in $C$.  [^6] Larger subgraphs encode more robust relationships, but mining $k$-node subgraphs for $k>5$ was intractable, and approximate subgraph mining tools were too imprecise for meaningful conclusions. For each $C$, we evaluate the percentage of the graphs in $\{S_{5}\}$ that are also subgraphs of $g$; i.e. we evaluate how well $g$ captures the most structurally salient musical relationships in $C$ (Table 6). On average, $70.33\%$ of each $\{S_{5}\}$ aligns perfectly with $g$, confirming each centroid captures the musically essential information of its corpus. To visualize these essential musical substructures, consider a 5-node subgraph common to the Mozart corpus and captured in its centroid (Figure 7). This reveals that a characterizing feature of the Mozart corpus is two consecutive major chords in the same major key, which in turn falls in a motif/pattern.
@@ 7 Conclusion and Future Work
Acknowledgments
This project is partially supported by the European Research Council under Europe’s Horizon 2020 program, grant #883313 (ERC REACH) .
@@ Appendix C Input Pieces for Section 6 Experiments
For details about the pieces used in the experiments, please consult our GitHub repository.777github.com/ilanashapiro/constraints_project Here, the 32 pieces used for the Structural Experiment (Section 6.1) can be found in the file: project/experiments/structural_distance/structural_distance_experiment/input_pieces.txt. The 210 set combinations of these pieces used in this experiment can be found in: project/experiments/structural_distance/structural_distance_experiment/set_combinations.txt. Finally, the pieces used for the Centroid Experiment (Section 6.2) can be found in the file: project/experiments/centroid/substructure_frequency_experiment/corpora/info.txt.

### v3

# Synthesizing Composite Hierarchical Structure from Symbolic Music Corpora

@@ Abstract
Western music is an innately hierarchical system of interacting levels of structure, from fine-grained melody to high-level form. In order to analyze music compositions holistically and at multiple granularities, we propose a unified, hierarchical meta-representation of musical structure called the structural temporal graph (STG). For a single piece, the STG is a data structure that defines a hierarchy of progressively finer structural musical features and the temporal relationships between them. We use the STG to enable a novel approach for deriving a representative structural summary of a music corpus, which we formalize as a nested NP-hard combinatorial optimization problem extending the Generalized Median Graph problem. Our approach first applies simulated annealing to develop a measure of structural distance between two music pieces rooted in graph isomorphism. Our approach then combines the formal guarantees of SMT solvers with nested simulated annealing over structural distances to produce a structurally sound, representative centroid STG for an entire corpus of STGs from individual pieces. To evaluate our approach, we conduct experiments verifying that structural distance accurately differentiates between music pieces, and that derived centroids accurately structurally characterize their corpora.
@@ 1 Introduction
A prevailing theory among Western music theorists and musicologists states that Western classical music exhibits an implicitly hierarchical structure [?]. While several different theoretical systems have been proposed to formalize this structural hierarchy [?], a widely accepted modern interpretation of the hierarchy states that melodies form the bottom, followed by harmonic contour, rhythmic patterns, disjoint and possibly overlapping motifs, and finally contiguous sections [???]. Together, this composite hierarchy encapsulates the overall structure of a piece.
To analyze musical structure computationally, many automatic approaches have been developed for extracting structure at single levels of the structural hierarchy [????], including methods to analyze sub-hierarchies within a single level [?]. However, music perception researchers have shown that the levels are not perceptually independent: they relate to one another in both “vertical” (structural) and “horizontal” (temporal) directions [?], interactions that have more recently been proven computationally  [?]. A comprehensive analysis of a piece thus must integrate these inter- and intra-level interactions into a unified model. Furthermore, while attempts have been made to generate structured music [???], to our knowledge, no existing research has examined generating music adhering to a complete, musically representative structural hierarchy, since no mechanism to computationally derive a complete structural specification from a desired music corpus currently exists. Such an encapsulation of music structure could also play a critical role in the generation of well-formed music by serving as a system of constraints on generative models.
@@
To address these challenges, we introduce the structural temporal graph (STG) as a unified model of complete musical structure. The STG is a $k$-partite directed acyclic graph whose levels form the structural hierarchy, and edges encode temporal relationships between adjacent levels. We use simulated annealing to develop a measure of structural distance between two STGs based on graph isomorphism, and to obtain the overarching structure of a corpus of pieces, we develop an approach to derive a representative centroid graph from a corpus of STGs. We formalize centroid derivation as a nested NP-hard combinatorial optimization problem extending the Generalized Median Graph problem [?], and propose a solution combining nested simulated annealing with the formal guarantees of SMT solvers to produce a structurally sound result. Our experiments show that structural distance accurately differentiates pieces, with its performance reliant on the complete hierarchy, and that derived centroids accurately structurally characterize their corpora. [^1] Paper code: https://github.com/ilanashapiro/stg_optimization
@@
    We formalize the music summarization problem as a nested NP-hard combinatorial optimization problem, and contribute a novel solution using both stochastic and SMT-based techniques.
@@ 3 Structural Temporal Graph
To address the lack of a fully automatic complete encapsulation of polyphonic musical structure, we introduce the structural temporal graph (STG), a unified meta-representation of musical structure that captures the levels of the music structural hierarchy and the temporal relationships between them. The STG is a $k$-partite directed acyclic graph (DAG), where each of the $k$ layers encodes a level in the music structural hierarchy. [^2] Individual levels themselves can form sub-hierarchies of increasing granularity, which the STG supports Following the modern music theoretic interpretation of the hierarchy [???], from top to bottom we denote the levels to be contiguous segmentation, motifs (both disjoint and overlapping), rhythmic contour, relative keys, functional harmonic chords, and melodic contour. The STGs we build include every level in the hierarchy except rhythmic contour, for which we were unable to access an analysis algorithm. We run individual analysis algorithms to generate each level of the hierarchy, which is elucidated in Section 6. Before formally defining the STG, we build intuition by walking through the derivation of an STG from an annotated piece.
@@ 4.1 Graph Alignment Annealing
GED is a NP-hard combinatorial optimization problem [?], intractable for most STGs. Most GED approximation algorithms are slow and of dubious accuracy, and more generalized than we require [?]. We thus introduce a new measure of structural distance computed with simulated annealing (SA), a stochastic optimization technique that estimates the global optimum of a discrete cost function. It comprises an objective “energy” function to minimize and a “move” function for generating a new solution from the current state. An annealer begins at a high temperature indicating the likelihood of accepting worse solutions to explore the solution space, and ends at a low temperature to refine near-optimal solutions [?].
@@ 5.1 Bi-Level Simulated Annealing
The centroid annealer’s energy function is given in Equation 5, where $C_{\text{aligned}}$ is the corpus aligned to the current centroid $A_{g}$, a process itself requiring SA as in Section 4.1 to obtain the optimal alignments. Finding the centroid $A_{g}$ minimizing Equation 4 is thus a nested NP-hard problem (GED and minimizing over these distances) requiring nested SA.
@@
As the centroid annealer’s temperature cools, the loss converges as the centroid is increasingly closely aligned to its corpus. Thus, as the centroid annealer’s temperature cools, we scale down the number of steps and max temperature of the nested alignment annealer. [^3] See Appendix A for precise cooling schedule
@@ 1. Every instance node must have exactly one prototype parent per feature
We first partition $A_{g}$ into pairs of consecutive instance levels without their prototypes (e.g. a segmentation/motif pair of instance levels), and optimize the instance constraints in Table 2 and relevant global constraints in Table 1 over each partition incrementally. We combine the results of each partition until we build a valid centroid subgraph of instance nodes. Then, we partition $A_{g}$ into single levels, each containing the instance nodes of that level and all prototype nodes for each instance feature at that level (e.g. segmentation instance nodes + section number prototype nodes). The instance constraints are already optimized; we need only optimize the prototype constraints in Table 2 and relevant global constraints in Table 1 over the possible prototypes. This gives us the complete, structurally sound centroid $A^{\prime}_{g}$ we seek. [^6] See Appendix B for centroid derivation visualization
@@ 6 Experiments
To evaluate our approach, we create a dataset of polyphonic, symbolic MIDI piano music from the Kunstderfuge and Classical Piano MIDI datasets. Since some single-level analyses we use to generate STGs require the data to be in audio and CSV format, we convert MIDI to CSV (with a manual script) and to MP3 (with Fluidsynth).
@@ 6.1 Structural Distance Evaluation
Musical Evaluation. To verify structural distance accurately differentiates pieces, we construct 210 STG sets from 32 pieces by J.S. Bach (21), Mozart (2), Beethoven (3), Schubert (2), and Chopin (4). Each set is a unique combination of 5 pieces, one from each composer, such that a piece’s duration is within 7 seconds of any other piece in the set, since structural distance between disparate length pieces could be due to STG size differences rather than the local structural variations we aim to distinguish. [^7] See Appendix C for input pieces/sets details We compute pairwise structural distances between the STGs in each set with our Section 4 graph alignment annealer (2000 iterations, max and min temperature of 2 and 0.01) on 8 Nvidia RTX 2080 GPUs with 11GB RAM. This results in 210 structural distance matrices, which we average to a single mean distance matrix.
We evaluate structural distance against three baselines over the same 210 piece combinations. Baseline 1 is the mean distance matrix obtained by taking the cosine similarity between feature vectors extracted from each MIDI file using Music21. Baseline 2 is the mean distance matrix over pairwise Stent weighted audio similarities (SWAS) for each piece’s paired MP3 file. SWAS is a composite audio similarity metric comprising zero-crossing rate, rhythm, chroma, spectral contrast, and perceptual similarity metrics, which we weigh equally. To demonstrate existing graph comparison metrics are insufficient, Baseline 3 is the Weisfeiler-Lehman (WL) Kernel applied pairwise to the STGs in each set, with five iterations and Neighborhood Hash (NH) as the base kernel. The WL Kernel iteratively refines node labels based on their neighbors, and uses a base kernel (in this case, NH, which efficiently captures local graph structure by distinguishing neighborhood configurations) to compare them [?]. Our ground-truth is the stylistic similarity indices between composers using human annotations and metadata from “The Classical Music Navigator” [?].
@@ 6.2 Centroid Evaluation
Mathematical Verification. As a correctness check, we evaluate derived centroids against constructed ground truth centroids. Given base STG $g$ with $|E|$ edges, we create a synthetic corpus $C_{k}$ of $k$ STGs by randomly adding $\lceil\frac{|E|}{2}\rceil$ valid edits to $g$, $k$ times, as in Section 6.1. By construction, $g$ must be a true centroid with min possible loss over $C_{k}$ (Equation 4). [^8] Proof in Appendix D We derive candidate centroid $g_{d}$ from $C_{k}$, and compare to $g$. We evaluate $g_{d}$ against the naive centroid $g_{n}$, i.e., the STG already in $C_{k}$ with min loss over the rest of $C_{k}$.
@@
Musical Evaluation. This experiment requires pieces with more length (and STG size) variation than Section 6.1, as centroids derived from very similar input graphs may be trivial by construction and fail to generalize to more diverse corpora. We thus relax the Section 6.1 relative duration restriction to 80 seconds, and generate centroid STGs for four corpora with pieces by Alkan (11), Chopin (8), Haydn (12), and Mozart (14) with the same architecture as before. [^9] See Appendix C for input pieces details
To verify each derived centroid $g$ musically characterizes its corpus $C$, we use rustworkx [?] to enumerate the set of all 5-node subgraphs ($\{S_{5}\}$) common to every STG in $C$, thus extracting the most structurally salient musical relationships in $C$ [^10] Larger subgraphs encode more robust relationships, but mining $k$-node subgraphs for $k>5$ was intractable, and approximate subgraph mining tools were too imprecise for meaningful conclusions For each $C$, we evaluate the percentage of the graphs in $\{S_{5}\}$ that are also subgraphs of $g$; i.e. we evaluate how well $g$ captures the most structurally salient musical relationships in $C$ (Table 6). On average, $70.33\%$ of each $\{S_{5}\}$ aligns perfectly with $g$, confirming each centroid captures the musically essential information of its corpus. To visualize these essential musical substructures, consider a 5-node subgraph common to the Mozart corpus and captured in its centroid (Figure 7). This reveals that a characterizing feature of the Mozart corpus is two consecutive major chords in the same major key, which in turn falls in a motif/pattern.
@@ Appendix C Input Pieces for Section 6 Experiments
For details about the pieces used in the experiments, please consult our GitHub repository. [^11] https://github.com/ilanashapiro/stg_optimization Here, the 32 pieces used for the Structural Experiment (Section 6.1) can be found in the file: project/experiments/structural_distance/structural_distance_experiment/input_pieces.txt. The 210 set combinations of these pieces used in this experiment can be found in: project/experiments/structural_distance/structural_distance_experiment/set_combinations.txt. Finally, the pieces used for the Centroid Experiment (Section 6.2) can be found in the file: project/experiments/centroid/substructure_frequency_experiment/corpora/info.txt.
@@ Appendix D Proof $g$ is True Centroid of Synthetic Corpus
Acknowledgments
This work is supported by National Science Foundation grants #1955457, #2220892, #2146151, and #2200333, and by the European Research Council under Europe’s Horizon 2020 program, grant #883313 (ERC REACH).